OpenStudy (anonymous):

can anyone help me with solving this application the yearly production of a 6 foot orange tree is 38 lbs and a 12 foot tree producing 84 lbs of oranges?

OpenStudy (anonymous):

Ah ok.

OpenStudy (anonymous):

So, what exactly is the question? It's just making a statement. Are we supposed to say how many pounds of fruit were created?

OpenStudy (anonymous):

no your given the numbers for each lbs of oranges two pairs what be (6,38) and (12,84)

OpenStudy (anonymous):

h represents height of a tree and p represents lbs of oranges

OpenStudy (anonymous):

this is the formula p-38=23/3(h-6)?

OpenStudy (anonymous):

Ah, ok! Let's see...

OpenStudy (anonymous):

Let's get slope first. y1-y2 / x1-x2 (6,38) and (12,84) So, 38-84 = -46 6-12 = -6 -23/-3 23/3 Liking your slope so far.

OpenStudy (anonymous):

y-y1 = m(x-x1) y-38 = 23/3(x-6) Yup! Now, we just need to simplify some more.

OpenStudy (anonymous):

were using p-p1=m(h-h1)

OpenStudy (anonymous):

y - 38 = (23/3)x - 6(23/3) y - 38 = (23/3)x - 138/3 y = 23/3x - 138/3 + 38 Now we have to try to combine the 38 with the -138/3. You tell me what we do here.

OpenStudy (anonymous):

substitute the values p1 and h1

OpenStudy (anonymous):

True. I was using the generic form of it. But we can go ahead and change the letters to specifics: p = 23/3h - 138/3 + 38 Now how do we add 138/3 and 38 together to make one fraction?

OpenStudy (anonymous):

p-38=23/3(h-6) are u familiar with this equation?

OpenStudy (anonymous):

Yup. I did say you were correct there. But recall what I said earlier: y - 38 = (23/3)x - 6(23/3) y - 38 = (23/3)x - 138/3 y = 23/3x - 138/3 + 38 Now we have to try to combine the 38 with the -138/3. You tell me what we do here.

OpenStudy (anonymous):

If you'd like, I'll go ahead and swap out the letters for you: p - 38 = (23/3)h - 6(23/3) p - 38 = (23/3)h - 138/3 p = 23/3h - 138/3 + 38 Now we have to try to combine the 38 with the -138/3. You tell me what we do here.

OpenStudy (anonymous):

the trouble im having is coming up with finding the numbers unknown for a certain lb of oranges for example 18 foot tree, 20 foot tree etc

OpenStudy (anonymous):

are we subtracting the 2 numbers ?

OpenStudy (anonymous):

are u there uber1337h

OpenStudy (anonymous):

Oh sorry, I was on facebook. lol. What you want to do is get a common divisor. We have -138/3 and 38. 38 as a fraction is 38/1, right?

OpenStudy (anonymous):

also u left off the h near the fraction

OpenStudy (anonymous):

always keep your variable near that fraction

OpenStudy (anonymous):

that last post was after you do the equation

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

that is the step after you set up your equation

OpenStudy (anonymous):

-138/3+ 38/1?

OpenStudy (anonymous):

p - 38 = (23/3)h - 6(23/3) p - 38 = (23/3)h - 138/3 p = 23/3h - 138/3 + 38 I don't see any missing H's :(

OpenStudy (anonymous):

-138/3h+38/1 oops

OpenStudy (anonymous):

Ah that. There is no h, though. That's why I was telling you we have to combine them.

OpenStudy (anonymous):

where did it go

OpenStudy (anonymous):

There never was one there to begin with. I'll show you the steps one at a time.

OpenStudy (anonymous):

it has to stay in the equation according to mymathlab

OpenStudy (anonymous):

p-38=23/3(h-6) We started here.

OpenStudy (anonymous):

Now, distribute the 23/3 to (h-6) We'd have (23/3 times h) - (23/3 times -6)

OpenStudy (anonymous):

Does that make sense so far?

OpenStudy (anonymous):

but we use the fraction in the equation which the h is representing the height of the orange trees so we have use the h

OpenStudy (anonymous):

And we do still use h. What do you get when you do this: "Now, distribute the 23/3 to (h-6) We'd have (23/3 times h) - (23/3 times -6) "

OpenStudy (anonymous):

let me show the equation in my math lab it says p-38=23/3h -46 was one of the answers to express p in terms of h remove parenthese

OpenStudy (anonymous):

And we will. Just do the equation I asked you to. :P

OpenStudy (anonymous):

does this make sense to you then u have to add 38 and simplify

OpenStudy (anonymous):

im not sure what you want me to do?

OpenStudy (anonymous):

Yup. That's what I've been asking you to do.

OpenStudy (anonymous):

oh -138/3?

OpenStudy (anonymous):

Yup. -138/3 plus 38.

OpenStudy (anonymous):

I've gotta go soon, so I'll tell you what we do:

OpenStudy (anonymous):

Since we know that 38 = 38/1 in fraction form, we multiply by 3/3 and get: (3(38))/1(3)

OpenStudy (anonymous):

im not sure how the got the 46? lol

OpenStudy (anonymous):

Hmm... 46 can't be right, as far as I can tell. lol

OpenStudy (anonymous):

ok but how do they get the 46 in the equation?

OpenStudy (anonymous):

Like I said, I think they made a typo. :(

OpenStudy (anonymous):

that was the first equation they did in the mathlab

OpenStudy (anonymous):

they were trying to find the number f lbs for a 18 foot tree and 20 foot tree

OpenStudy (anonymous):

p - 38 = (23/3)h - 6(23/3) p - 38 = (23/3)h - 138/3 p = 23/3h - 138/3 + 38 So now we have this: p = 23/3h - 138/3 + 414/3 By adding the 414/3 and -138/3, we have: 276/3

OpenStudy (anonymous):

well its no typo its in my study plan module i plug in the wrong answer and they gave their answer

OpenStudy (anonymous):

there is about 4 parts to this word problem application

OpenStudy (anonymous):

remember its a word problem u are find number of lbs per 18 foot tree and and 20 foot tree here

OpenStudy (anonymous):

were using the same equation in the first half to answer these problems

OpenStudy (anonymous):

Alright, so let's do this one step at a time. I see what the problem is.

OpenStudy (anonymous):

Remember this equation? p-38=23/3(h-6)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome. First of all, let me double check. Did you get that by yourself? And if not, are you sure you know how we got to that equation from just the two points that were given to us?

OpenStudy (anonymous):

Still here? Gotta go soon.

OpenStudy (anonymous):

Lemme know if you still want to do this. I need to head to bed soon.

OpenStudy (anonymous):

hello u there

OpenStudy (anonymous):

yes im in my study plan online in classroon portal

OpenStudy (anonymous):

Sigh... I guess since you at least tried a little, I'll tell you what it is... p = (23/3)h - 8 I wanted to explain it more in depth, but you kinda left abruptly. Just hit me up some other day and I'll try to explain it more in depth. But basically, this should explain it: p - 38 = (23/3)h - 6(23/3) p - 38 = (23/3)h - 138/3 p = (23/3)h - 138/3 + 38 p = (23/3h) - 46 + 38 p = (23/3)h - 8 Anyhow, the reason I thought -46 was a typo is because I didn't realize that -138/3 = -46. Remember to ask me in the future for more information if you see me.

OpenStudy (anonymous):

Ah, good to see you're back then. :P How do you like that answer?

OpenStudy (anonymous):

the computer will plug in the answers when u give the wrong answers but some steps are skipped

OpenStudy (anonymous):

The -46 came from the -138/3 that we got earlier. So just use this work and it should be all happy.

OpenStudy (anonymous):

im here waiting

OpenStudy (anonymous):

For what?